# Distance Recomputator and Topology Reconstructor for Graph Neural Networks

### v1

@@ Abstract
Graph Neural Networks (GNNs) have gained prominence in semi-supervised learning for graph representation due to their ability to capture intricate node relationships. Recently, there is a trend for k-hop structure learning for GNNs. While GAMLP [ZYS+22] trains an MLP layer for each k-hop domain, ImprovingTE [YWZL23] enhances this approach by injecting contextualized substructure information to effectively utilize the k-hop structure. However, those traditional k-hop sampling approaches have largely relied on sampling performance, which limits the upper bound of accuracy and made the outcome unstable. To address this limitation, inspired by ”coraset selection”[GZB22] idea, we develop a novel approach that facilitates k-hop structure sampling and message passing, extending the reach and depth of information flow within the graph. To tackle the challenges of mislabeling and inaccuracies in datasets, we introduce two innovative models: the ”Distance Recomputator” and the ”Topology Reconstructor.” The Distance Recomputator recalibrates the distances between nodes, thereby refining node representations and interactions in a more accurate and context-aware manner. Complementing this, the Topology Reconstructor dynamically adjusts local graph structures, enhancing the model’s adaptability to complex and evolving graph topologies. Our experimental results indicate significant performance enhancements of these models over existing benchmarks.
1University of Wisconsin-Madison,
2University of Notre Dame
@@ 1 Introduction
Graph Neural Networks (GNNs) have emerged as a powerful tool in the realm of machine learning, adept at capturing the complex relationships inherent in graph-structured data. From social network analysis to molecular structure interpretation, GNNs have demonstrated remarkable versatility. However, their ability to model dependencies and interactions within graph structures is fundamentally constrained by the methods used to compute and represent node relationships. Besides, GNN in graph local structure learning (K-hop) has became more and more popular, such as [WYHL21]. Traditional GNN architectures, while effective, often struggle with efficiently encoding the dynamic and intricate topologies of real-world graphs. This limitation motivates the need for more adaptive and robust models that can better capture the nuances of graph data.
Current GNN models predominantly rely on static node representations and fixed neighborhood aggregation schemes, leading to suboptimal performance in scenarios where graph topology is not only complex but also dynamic. Additionally, the computation of node distances in large graphs often incurs significant computational overhead, thereby limiting scalability. These challenges are accentuated in applications involving large-scale graphs or rapidly evolving network structures, such as in communication networks or dynamic social graphs. The limitations in handling varying node distances and adapting to topological changes prompt the exploration of more flexible and efficient approaches.
To address these challenges, we introduce two novel models: the ”Distance Recomputator” and the ”Topology Reconstructor.” The Distance Recomputator is designed to efficiently recalibrate node distances within a specified k-hop domain, leveraging a dynamic encoding scheme that adapts to changes in node proximity and graph density. This allows for a more nuanced representation of node relationships, enhancing the accuracy of dependency modeling in the network. Complementing this, the Topology Reconstructor dynamically adjusts the local network topology, enabling the model to respond to structural changes in the graph. By integrating these models into standard GNN frameworks, we propose a solution that not only addresses the static nature of traditional models but also introduces a level of adaptability hitherto unseen in GNN architectures.
Our experimental evaluation of these models demonstrates a marked improvement in performance across a range of benchmark datasets, particularly in tasks involving dynamic or large-scale graphs. The Distance Recomputator shows enhanced efficiency in recalculating node distances, leading to more accurate node representations and predictions. Similarly, the Topology Reconstructor proves effective in adapting to topological changes, thereby maintaining model robustness. Complementing these empirical results, we provide a comprehensive theoretical analysis that elucidates the mechanisms by which these models achieve superior performance. This analysis not only validates our experimental findings but also contributes to a deeper understanding of the underlying principles governing effective GNN design.
@@ 2.1 Topology imbalance Problem
”Not every link are useful as their topology reflected”. The topology of a graph in GNNs does not always accurately reflect the significance of each connection. [LLC+23] Consider a social network: an account labeled ”math” might follow one food-maker and ten mathematical educators. In this context, the link to the food-maker is less relevant for label prediction tasks. Current graph neural models, such as GNNs, indiscriminately propagate, aggregate, and update information from all 1-hop neighbors, including less relevant connections like the food-maker. This issue highlights the ’Topology Imbalance Problem,’ where certain links (e.g., those to math educators) are more informative and relevant than others (e.g., the food-maker link).
Although models like Graph Attention Networks (GAT) assign an importance score to each link, they rely on a synchronous aggregator, treating all nodes within the same hop identically during propagation and update phases. To address this imbalance, our paper introduces the ’Distance Recomputator’ model, which recalculates node distances to better reflect the relevance of links. Complementarily, we propose an ’Asynchronous Aggregator’ that enables nodes to be aggregated based on these recalculated distances, allowing for more nuanced and context-sensitive information processing.
2.2 The Shortcoming of Synchronous Aggregator
Synchronous aggregators in Graph Neural Network process neighbors in one aggregation, the synchronous approach can inadvertently amplify the impact of less relevant or erroneous links, potentially degrading the model’s performance. It fails to differentiate between the varying levels of relevance among neighboring nodes, treating all connections as equally significant during the aggregation process. Our proposed asynchronous aggregator aims to mitigate this shortcoming by allowing for selective, relevance-based aggregation of neighborhood information, thereby enhancing the model’s accuracy and robustness against irrelevant or misleading connections.
Some current work shows the asynchronous processing of GNNs, such as AEGNN[SGS22], which designs update rules that restrict recomputation of network activations only to the nodes based on each new event, while Gated Graph Sequence Neural Networks [LTBZ17] deploys a method that can use gated recurrent units to extend to output sequences of GNNs to realize asynchronous processing. While all above methods give solution for asynchronous processing in message passing, they did not give the solution at a topology level.
@@ 2.3 Heat Diffusion in Graph Neural Networks
Heat Diffusion in GNNs is inspired by the physical phenomenon of heat transimisision, where heat (information) dissipates as it moves away from the heat source.[TDKF16] Our work also explores the concept of heat diffusion in the context of GNNs, particularly in k-hop message passing. We simulate thermodynamic properties, where information intensity decreases progressively during propagation. By applying this concept to GNNs, we introduce a novel method to control information flow, ensuring that the propagation of data mimics natural attenuation over distance. This method not only provides a more realistic approach to information dissemination in networks but also helps in reducing the noise and enhancing the signal-to-noise ratio in the message-passing process. More details about this innovative approach can be found in our supplementary document, ”Heat Diffusion in GNNs”.
More details in this link: Heat Diffusion in GNNs
@@ 2.4 Paper Contribution
  • K-hop Diffusion Message Passing Strategy: We develop a message passing framework enabling vertices to learn k-hop information at each propagation stage. Post-propagation, node representations and inter-node distances are recalibrated, reflecting the acquired k-hop information. This layer-by-layer information dissemination, incorporating heat diffusion within the k-hop domain, refines node characteristics and relational metrics continuously.
  • Distance Recomputator: Our model employs an attention mechanism to recalculate node distances, considering both k-hop topology and vertex features. This recalibration allows for accurate, context-aware node relationship representations within the graph.
  • Topology Reconstructor: Introducing the first k-hop topology reconstruction method, our model leverages k-hop topology from sampling to perform reconstruction based on computed ”similarity distances.” Nodes exceeding a similarity distance threshold are repositioned to optimize network configurations for enhanced learning outcomes.
3 Related Work and Innovation
@@ 3.1 K-hop Message Passing
The prevailing K-Hop models in graph neural networks predominantly emphasize K-time message passing with 1-hop propagation schemes. Existing models lack the capability for k-hop message passing with k-hop propagation in a single iteration, or for one-time k-hop sampling methods. For instance, the enhancing multi-hop connectivity model reevaluates connectivity in the k-hop neighborhood by sampling k-hop neighbors and reweighting high-order connections, rewarding highly-related nodes and penalizing less-correlated ones [LJZ+22]. Another model, KP-GNN, incorporates peripheral embeddings to enrich representation learning at each layer [FCL+22]. Our model breaks new ground by efficiently sampling k-hop neighborhood information and realizing k-hop message passing within this domain. By exploring k-hop neighborhoods in each propagation cycle, our framework captures more locality information than conventional 1-hop methods, thereby enhancing label prediction accuracy for the central node with a comprehensive k-hop perspective.
@@ 3.2 Diffusion Propagation in Graph Neural Networks
[…]
3.3 Graph Imbalancing Problems
The issue of label and topology imbalance in graph neural networks has been relatively underexplored. Zhao et al. [ZLZW22] proposed adjusting edge weights to mitigate topology imbalance effects without altering the graph structure. In contrast, our research introduces two innovative models to address this challenge comprehensively. The first, GKHDDRA, utilizes hop jumping to adjust the graph’s topology. The second, GDRA, fine-tunes the dataset by discarding irrelevant edges and forming new connections between strongly similar nodes, reshaping the graph’s topology to achieve a more balanced and representative structure.
4 Basic Methodology on Graph Neural Networks
Consider an undirected graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, with the node set $\mathcal{V}$ and edge set $\mathcal{E}$. The adjancy matrix A $\in$ $R^{N\times N}$ describes the existence of connections among nodes. Every node $v_{i}\in\mathcal{V}$ has an associated feature vector $x_{i}\in R^{d\times 1}$, so the whole feature vector space can be represented as $X=[x_{1},x_{2},...,x_{N}]^{T}$, so that the goal is to predict the labels of the remaining nodes. For a semi-supervised node classification task on a graph, the labels $Y_{L}$ are only available for a subset of the noes (training set), so the goal is to predict the labels of the remaining nodes.
@@ 4 Basic Methodology on Graph Neural Networks
where $h_{v}^{(k)}$ is the feature vector of node v at the $k^{th}$ layer. We initialize the $h_{v}^{(0)}=x_{v}$, and the $\mathcal{N(v)}$ is a set of nodes connected to $v$. Existing work are focus on how to improve the work within 1-hop $Aggregate^{(k)}(\dot{)}$ and $Combine^{(k)}(\dot{)}$ operations. For example, GCN employs a convolution operation with the following rule
$H^{(l+1)}=\sigma(\hat{A}H^{(l)},\theta^{(l)})$ (2)
where $H^{(l)}$ are the output features in the $l^{th}$ layer. We initialize the $H^{(0)}=X$, $\hat{A}=\hat{D}^{-\frac{1}{2}}(A+I)\hat{D}^{-\frac{1}{2}}$ is the Laplacian normalized adjacency matrix, and the $\theta^{(l)}$ is the $l^{th}$ layer weights of the neural network. GNN models stack GNN layer to explore the representation in larger domain, take two-layer GCN model as an example, the represetnation can be calculated as $Z=softmax(\hat{A}\sigma(\hat{A}X\theta^{(0)})\theta^{(1)})$. As for node classification task with the labeled nodes corresponding to $Y_{L}$, the objective function is
@@
4.1 The In-Depth Learning of Graph Neural Networks
Contemporary Graph Neural Network (GNN) models predominantly adopt a 1-hop depth message passing paradigm, where each propagation cycle explores information at a singular depth level. In contrast to GraphSAGE [HYL17], which samples along k-depth but aggregates at every depth, our approach aims to develop a model capable of both sampling and aggregating in a k-hop domain. This method is anticipated to yield a more comprehensive understanding of the local structural context within the network.
We propose an in-depth learning algorithm for k-hop sampling and message passing in GNNs, characterized by the following steps:
In the proposed Graph Neural Network model, we initiate with a Static Preprocessing Step, where a substantial number of neighbors for each node are sampled. This phase involves constructing a k-hop adjacency matrix for every node, effectively capturing and storing the k-hop neighborhood information. This process is computationally efficient, with a complexity of $\mathcal{O}(\mathcal{V*K})$, where $\mathcal{V}$ represents the total number of vertices in the graph. Following this, the model enters the Dynamic Update Phase during propagation and update cycles. In this phase, a smaller, more focused subset of k-hop neighbors for each node is dynamically sampled. This targeted approach allows for the refinement of the pre-processed k-hop neighborhood matrix, ensuring that the model remains up-to-date and relevant in representing the evolving graph structure. The computational complexity of this phase is maintained at $\mathcal{O(K*R)}$, with $\mathcal{R}$ indicating the limited number of vertices sampled in the k-hop domain during each propagation step. Together, these steps create a robust framework for in-depth learning in Graph Neural Networks, optimizing both the breadth and depth of neighborhood information processing.
The k-hop attentive message passing strategy is underpinned by both dynamic and static sampling paths. This mechanism is grounded in the principles of graph diffusion networks, while the computation of distance reevaluation indicators is inspired by and derived from the architecture of the Graph Attention Network (GAT) [VCC+18]. This novel approach aims to harness the strengths of both static and dynamic sampling methodologies to enhance the depth and accuracy of message passing in GNNs.
Input :  Graph $\mathcal{G(V,E)}$; depth $K$; Sampling number per hop $N$; Input features: $Z$; Adjacency Matrix: A;
Output : NGH //K-hop Sampling Storage, a 3D dictionary
NGH[0][:] $\leftarrow$ A

### v2

# Adaptive Reconstruction for Graph Neural Networks

@@ Abstract
Graph Neural Networks (GNNs) have become fundamental in semi-supervised learning for graph representation, leveraging their ability to capture complex node relationships. A recent trend in GNN research focuses on adaptive k-hop structure learning, moving beyond fixed-hop aggregation to more flexible and dynamic neighborhood selection. While GAMLP Zhang et al. (2022) employs separate MLP layers for each k-hop domain and ImprovingTE Yao et al. (2023) enhances this by injecting contextualized substructure information, these methods still rely heavily on predefined sampling strategies, which may limit their ability to generalize and maintain stable accuracy. To address these limitations, we propose an adaptive reconstruction framework that dynamically refines k-hop structure learning. Inspired by ”coreset selection” Guo et al. (2022), our approach adaptively reconstructs node neighborhoods to optimize message passing, ensuring more effective and context-aware information flow across the graph. To further enhance structural robustness, we introduce two key modules: the Distance Recomputator and the Topology Reconstructor (DRTR). The Distance Recomputator reassesses and recalibrates node distances based on adaptive graph properties, leading to improved node embeddings that better reflect latent relationships. Meanwhile, the Topology Reconstructor dynamically refines local graph structures, enabling the model to adapt to evolving graph topologies and mitigate the impact of noise and mislabeled data. Empirical evaluations demonstrate that our adaptive reconstruction framework achieves significant improvements over existing k-hop-based models, providing more stable and accurate performance in various graph learning benchmarks.
@@ 1 Introduction
Existing GNN architectures predominantly adopt static node representations and predefined aggregation schemes, which can be suboptimal when handling graphs with dynamic topologies. Furthermore, the computation of node distances in large-scale graphs introduces significant overhead, making it challenging to scale these models effectively Kung et al. (2024); Wang et al. (2024). These limitations highlight the need for a more flexible and efficient approach that can dynamically reconstruct graph structures, allowing models to better adapt to changes in connectivity patterns and improve representation quality.
To address these challenges, we propose an adaptive reconstruction framework that refines both node distance calculations and local topological structures to enhance message passing efficiency. This framework introduces two key components: the Distance Recomputator and the Topology Reconstructor (DRTR). The Distance Recomputator dynamically recalibrates node distances within the k-hop domain by incorporating an adaptive encoding mechanism that adjusts to changes in graph density and node proximity. By refining distance computations, it provides a more precise representation of node relationships, which improves downstream predictions and model stability. The Topology Reconstructor complements this by continuously adjusting local graph connectivity, ensuring that the model remains robust in scenarios where structural variations occur. This adaptive mechanism allows GNNs to dynamically reconstruct neighborhood structures, making them more resilient to graph shifts and variations in node interactions.
Through extensive experiments on multiple benchmark datasets, we demonstrate that our adaptive reconstruction approach significantly improves performance in tasks involving both static and dynamic graphs. The Distance Recomputator proves effective in optimizing node representations, reducing the impact of noisy or misleading edges while enhancing the quality of learned embeddings. Meanwhile, the Topology Reconstructor enables models to better align with evolving graph structures, leading to more stable and accurate predictions. In addition to empirical validation, we provide a theoretical analysis that explains the benefits of reconstruction in GNN learning dynamics, offering deeper insights into why adaptively refining graph structures leads to superior model generalization.
@@ 2.1 Topology Imbalance Problem
The structural connections in a graph do not always accurately reflect the significance of each link. As stated in Liu et al. (2023), ”not every link is as useful as its topology suggests.” Consider a social network where an account labeled ”math” follows ten mathematical educators and one food-maker. For tasks such as label prediction, the connection to the food-maker contributes little relevant information, yet traditional Graph Neural Networks (GNNs) aggregate information from all 1-hop neighbors indiscriminately. This results in a phenomenon we define as the Topology Imbalance Problem, where some connections (e.g., links to math educators) carry significantly more useful information than others (e.g., the food-maker link).
While models such as Graph Attention Networks (GAT) assign importance scores to each link, they still rely on synchronous aggregation, where all nodes within the same hop are processed simultaneously. This limitation prevents the model from prioritizing more relevant connections over less meaningful ones during message passing. To overcome this issue, we propose the Distance Recomputator, which dynamically recalculates node distances to better reflect the contextual importance of each link. Additionally, we introduce the Asynchronous Aggregator, which enables nodes to be aggregated based on these recomputed distances, allowing for a more adaptive and information-sensitive update mechanism.
2.2 Limitations of Synchronous Aggregation
Most GNN architectures employ synchronous aggregators, where information from a node’s immediate neighbors is aggregated simultaneously. For example, Graph Convolutional Networks (GCN) aggregate neighbor information by weighting it with the inverse node degree, while GAT uses an adjacency matrix mask and attention scores to prioritize neighbors. Similarly, GraphSAGE samples a subset of neighbors before applying an aggregation function. Despite their differences in implementation, these models share a common characteristic: all nodes in the same layer are processed concurrently.
A major drawback of synchronous aggregation is its inability to differentiate between links of varying importance in a timely manner. Since all neighboring nodes are updated simultaneously, irrelevant or even erroneous connections can propagate misleading information, reducing model robustness. This issue becomes more pronounced in large and complex graphs, where spurious or noisy edges can significantly degrade performance.
To address this limitation, we propose an asynchronous aggregation mechanism that selectively integrates information from different neighbors based on their relevance. Unlike traditional synchronous approaches, our method dynamically determines the order and priority of aggregation, reducing the negative influence of uninformative links while preserving critical structural relationships. Some recent works have explored asynchronous processing in GNNs, such as AEGNN Schaefer et al. (2022), which updates node activations only upon new events, and Gated Graph Sequence Neural Networks Li et al. (2017), which employ gated recurrent units to process sequences asynchronously. However, these methods primarily focus on message passing rather than addressing topological reconstruction at a structural level. Our approach directly enhances the representation of graph topology, offering a more comprehensive solution.
@@ 2.3 Heat Diffusion in Graph Neural Networks
The concept of heat diffusion offers a valuable perspective for understanding information propagation in Graph Neural Networks (GNNs). Drawing inspiration from thermodynamics, where heat gradually dissipates as it moves away from its source Thanou et al. (2016), we analyze how information in GNNs follows a similar pattern of attenuation during multi-hop message passing. As information propagates across a graph, its influence weakens with increasing distance from the source node, often leading to a decline in signal quality and an accumulation of noise. Traditional GNN architectures typically assume uniform propagation across all hops, failing to account for this natural decay in information relevance.
To address this limitation, we introduce a distance-aware reconstruction framework that models message passing through a heat diffusion-inspired mechanism. Our approach regulates the transmission strength based on a distance-dependent attenuation function, ensuring that information decays in a controlled and meaningful manner. By integrating this adaptive reconstruction process, we prevent the excessive spread of weak or irrelevant signals while reinforcing the retention of important structural patterns. This refinement enhances both the signal-to-noise ratio and the interpretability of learned node representations, making the model more robust to graph sparsity and noise.
By incorporating heat diffusion principles into the reconstruction process, our method achieves a more balanced and adaptive message propagation strategy. This approach not only improves the efficiency of information flow but also mitigates common issues such as oversmoothing and excessive feature mixing in deep GNN models. Additional theoretical analysis and mathematical formulations of our heat diffusion-based reconstruction mechanism are provided in our supplementary document, ”Heat Diffusion in GNNs.”
@@ 2.4 Paper Contribution
  • K-hop Diffusion Message Passing: We propose a message passing framework that integrates k-hop information at each propagation step. After each update, the model recalibrates node embeddings and distances to reflect newly acquired information. A heat diffusion-inspired mechanism controls attenuation, preserving meaningful connections while reducing noise.
  • Distance Recomputator: Our model recalculates node distances using an adaptive attention mechanism that considers both k-hop topology and feature interactions. It continuously adjusts distances based on contextual relevance, improving message passing and filtering irrelevant edges.
  • Topology Reconstructor: We introduce a k-hop topology reconstruction method. our model leverages k-hop topology from sampling to perform reconstruction based on computed ”similarity distances.” Nodes exceeding a similarity distance threshold are repositioned to optimize network configurations for enhanced learning outcomes.
3 Core Ideas of DRTR
@@ 3.1 K-hop Message Passing
Most existing k-hop models in graph neural networks rely on multiple rounds of 1-hop message passing to capture higher-order information. However, they lack the ability to perform k-hop message passing with k-hop propagation in a single iteration or efficiently sample k-hop neighborhoods in one step. For example, the enhancing multi-hop connectivity model refines k-hop neighborhoods by sampling k-hop neighbors and reweighting high-order connections, strengthening highly related nodes while down-weighting weaker ones Liu et al. (2022). Similarly, KP-GNN enriches node representations by incorporating peripheral embeddings at each layer Feng et al. (2022).
Our model introduces an efficient k-hop sampling strategy that enables direct k-hop message passing, capturing local structures more effectively than traditional 1-hop methods. By integrating k-hop information in each propagation step, our approach provides a more comprehensive graph perspective, leading to improved label prediction accuracy for central nodes.
@@ 3.2 Diffusion Propagation in Graph Neural Networks
We propose a diffusion-based propagation model that builds upon traditional approaches such as the Diffusion Decent Network. One well-known model, MAGNA Wang et al. (2021), propagates information by updating vertex and edge embeddings while recalculating distances for both one-hop and multi-hop neighbors, considering all paths up to $k$ hops. Another approach, ”Diffusion Improves Graph Learning” Klicpera et al. (2019), approximates the diffusion equation using an infinite series, improving computational efficiency on large-scale graphs.
Our model extends these concepts by simulating heat diffusion within the k-hop domain. This diffusion-inspired framework regulates message passing, dynamically adjusting the flow of information to reduce noise accumulation and enhance representation quality. By incorporating heat diffusion into k-hop propagation, our model provides a structured and adaptive approach to information distribution in graph neural networks.
3.3 Graph Imbalance Problems
Label and topology imbalance in graph neural networks remain significant but underexplored challenges. Zhao et al. Zhao et al. (2022) addressed topology imbalance by adjusting edge weights without modifying the graph structure.
We introduce two models that actively reconstruct graph topology to improve balance. The first, GKHDDRA, applies hop jumping to restructure the graph and refine connectivity. The second, GDRA, enhances the dataset by removing weak edges and forming new connections between strongly similar nodes. These adjustments optimize graph structure, leading to a more balanced and representative network, ultimately improving learning performance.
4 Architectures of DRTR
The Distance Recomputator and Topology Reconstructor (DRTR) is designed to enhance Graph Neural Networks (GNNs) by addressing two major challenges: (1) capturing long-range dependencies while maintaining computational efficiency, and (2) dynamically adjusting the graph topology based on learned representations. The K-hop Diffusion Attention Layer, a core component of DRTR, integrates static preprocessing, multi-hop diffusion, and dynamic topology updates to optimize message passing and representation learning.
Graph-based learning problems often rely on an underlying undirected graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, with the node set $\mathcal{V}$ and edge set $\mathcal{E}$. The adjancy matrix A $\in$ $R^{N\times N}$ describes the existence of connections among nodes. Every node $v_{i}\in\mathcal{V}$ has an associated feature vector $x_{i}\in R^{d\times 1}$, so the whole feature vector space can be represented as $X=[x_{1},x_{2},...,x_{N}]^{T}$, so that the goal is to predict the labels of the remaining nodes. For a semi-supervised node classification task on a graph, the labels $Y_{L}$ are only available for a subset of the noes (training set), so the goal is to predict the labels of the remaining nodes.
@@ 4 Architectures of DRTR
where $h_{v}^{(k)}$ represents the feature vector of node $v$ at the $k^{th}$ layer. The initialization follows $h_{v}^{(0)}=x_{v}$, where $\mathcal{N}(v)$ denotes the set of neighboring nodes connected to $v$.
Existing works primarily focus on improving the 1-hop aggregation $\mathcal{AGGREGATE}^{(k)}(\cdot)$ and combination $\mathcal{COMBINE}^{(k)}(\cdot)$ operations. For example, Graph Convolutional Networks (GCN) utilize a convolution operation defined as:
$H^{(l+1)}=\sigma(\hat{A}H^{(l)}\Theta^{(l)})$ (2)
where $H^{(l)}$ represents the feature matrix at the $l^{th}$ layer, initialized as $H^{(0)}=X$. The term $\hat{A}=\hat{D}^{-\frac{1}{2}}(A+I)\hat{D}^{-\frac{1}{2}}$ is the Laplacian-normalized adjacency matrix, and $\Theta^{(l)}$ is the learnable weight matrix at the $l^{th}$ layer.
GNN models stack multiple layers to capture hierarchical representations over a larger neighborhood. Taking a two-layer GCN as an example, the final node representation is computed as:
$Z=\text{softmax}(\hat{A}\sigma(\hat{A}X\Theta^{(0)})\Theta^{(1)})$ (3)
For a node classification task, given labeled nodes corresponding to $Y_{L}$, the objective function is defined as:
@@
4.1 In-Depth Learning in Graph Neural Networks
Contemporary GNNs predominantly employ a 1-hop message passing paradigm, where each propagation step explores information within a single depth level. Unlike GraphSAGE Hamilton et al. (2017), which samples along a $k$-depth domain but aggregates at every depth, our approach integrates both sampling and aggregation within a $k$-hop neighborhood. This enables a more comprehensive understanding of local structural dependencies within the graph.
We propose an in-depth learning algorithm for $k$-hop sampling and message passing in GNNs, comprising the following key components:
1. Static Preprocessing Step: A large number of neighbors are pre-sampled for each node, constructing a $k$-hop adjacency matrix that encapsulates extended neighborhood information. This process achieves efficient storage of structural dependencies with a computational complexity of $\mathcal{O}(\mathcal{V}K)$, where $\mathcal{V}$ denotes the total number of vertices.
2. Dynamic Update Phase: During propagation, a refined subset of $k$-hop neighbors is dynamically sampled, updating the precomputed neighborhood structure. This step ensures that the model adapts to evolving graph topologies while maintaining computational efficiency at $\mathcal{O}(KR)$, where $\mathcal{R}$ represents the number of sampled vertices per propagation step.
The $k$-hop attentive message passing strategy leverages both static and dynamic sampling pathways. This mechanism is inspired by graph diffusion networks and integrates distance reevaluation indicators derived from the Graph Attention Network (GAT) Veličković et al. (2018). By synthesizing static and dynamic sampling methodologies, our approach enhances both the depth and accuracy of message passing in GNNs.
Input :  Graph $\mathcal{G(V,E)}$, Depth $K$, Sampling number per hop $N$,
Input features $\mathbf{Z}$, Adjacency Matrix $\mathbf{A}$
Output : $\mathcal{NGH}$: K-hop Sampling Storage (3D dictionary)
$\mathcal{NGH}^{(0)}\leftarrow\mathbf{A}$
  // Initialize neighborhood